# Spectral methods for wedge and corner flows: The Fourier–Kontorovich–Lebedev integral transform

### 11 Mar 2026 v1

@@ I.1 The origin of the Kontorovich–Lebedev Transform
The KL integral transform is a relatively less known yet powerful integral transform, particularly well suited for solving boundary-value problems in wedge-shaped or conical geometries. The method was first developed in the late 1930s by the Russian mathematicians Mikhail Iosifovich Kontorovich (1906–1987) and Nikolai Nikolaevich Lebedev (1911–1994), who were affiliated with the Leningrad Physico-Technical Institute of the Academy of Sciences of the U.S.S.R., now known as the A. F. Ioffe Institute in Saint Petersburg. Using this approach, boundary-value problems in diffraction theory [7, 8] and electrodynamics [9] were solved by Kontorovich and Lebedev. The foundational mathematical theory of the integral transform was subsequently developed by Lebedev [10, *lebedev1946french, 12, 13]. A concise textbook presenting detailed solutions to applied mathematics problems—including applications of KL transforms—was coauthored by Lebedev together with Skal’skaia and Uflyand, first published in Russian and later translated and edited by Silverman [14, Ch. 6, p. 194].
In contrast to more familiar transforms, the KL transform is classified as an index integral transform [15, 16, 6]. These form a class of integral operators in which the transformed function depends on an index parameter, typically linked to special functions [17]. Over recent years, numerous significant contributions to the foundations of the KL transform have been made by Yakubovich and his collaborators [18, 19, 20, 21, 22]. They advanced the theory of the KL transform by developing related integral and convolution operators and establishing connections with other integral transforms. Together, these developments established the KL transform as a versatile and rigorous tool for tackling complex boundary-value problems in mathematical physics. For a compact overview of this outstanding integral transform, including tables of transforms for common special functions, see the classic textbook by Erdélyi et al. [1, Ch. XII, p. 173].
@@ I.3 Applications in linear elasticity
Unlike its extensive use in electromagnetic and acoustic diffraction, the KL transform has seen comparatively limited application in linear elasticity [47]. The integral transform has been used to derive analytical solutions of the Navier–Cauchy equations for wedge-like geometries, providing a powerful tool for addressing stress and displacement fields in solids with angular boundaries. Early applications by Kassir and collaborators [48, 49] focused on solving crack problems using Papkovich–Neuber representations, yielding the stress distribution within the solid. To understand crack propagation and stress concentration in layered wedge-shaped structures, a periodic crack problem in an elastic layered wedge was analyzed by Pozharskii et al. [50]. In a related study, strip and elliptical cuts located in the middle of an elastic wedge were investigated by Aleksandrov and Pozharskii [51].
@@ I.4 Applications in low-Reynolds-number flows
The study of low-Reynolds-number flows near three-dimensional corners has provided fundamental insights into viscous fluid behavior in angular confinements. Stokes flows in wedge geometries were first studied by Hasimoto and collaborators, who analyzed a sequence of wedge geometries, starting with a right-angled wedge [58], followed by obtuse wedges [59, 60, 61, 62], and semi-infinite planes [63]. The case of salient wedges was later investigated by Kim [64]. In these problems, the solution is obtained using a Fourier transform along the axial direction and a KL transform along the radial direction. These studies enabled the calculation of hydrodynamic mobility functions of colloidal particles in the far-field limit. Remarkably, it has been shown that when the wedge opening angle falls below a threshold value, an infinite sequence of eddy rings forms around the direction of the point force [65]. The Stokes flow in a semi-infinite wedge was later investigated by Shankar using the method of vector eigenfunctions [66].
@@
Although the KL transform was historically developed in wave physics, the present review concentrates on its role in viscous flows near wedges and corners. We revisit key techniques introduced in earlier works and summarize the solution methodology as adapted to incompressible Stokes flow. We focus on the steady Stokes equations, so that time does not explicitly appear in the formulation, allowing the analysis to concentrate on spatial variations of the flow within wedge geometries. Since our approach combines Fourier transforms along the axial direction with the KL transform, we refer to this as the Fourier–Kontorovich–Lebedev (FKL) transform, representing a synthesis of both techniques.
Our discussion is restricted to systems in which the fluid motion is driven by either a point force or a point torque, and we focus on the physically relevant case of no-slip boundary conditions imposed on the wedge surfaces. Other types of boundary conditions can also be analyzed using the same methodological framework presented in this article. The goal of this review is to progressively develop the theoretical approach underlying the FKL transform, clarify its practical implementation for wedge geometries, and consolidate existing results in order to provide a coherent foundation for future advances in corner and wedge hydrodynamics. This review serves as an introductory read for those wishing to explore the modeling of flow problems near wedges and corners, progressively presenting the underlying mathematical framework and methods.
The remainder of the article is organized as follows. In Sec. II, we present the equations governing fluid motion and outline the framework used to obtain solutions based on the Papkovich–Neuber representation. Section III introduces the FKL transform, first by reviewing the relevant properties, and then by representing the solution form in FKL space. In Sec. IV, we summarize previously known solutions for a point force and point torque acting in a wedge-shaped geometry for arbitrary orientations. Finally, Sec. V provides a discussion and concluding remarks.
@@ II Governing equations
The system under investigation is three dimensional and consists of a viscous, incompressible fluid confined within a wedge-shaped domain, bounded by two walls located at $\theta=\pm\alpha$, where the wedge semi-opening angle $\alpha\in(0,\pi/2]$; see Fig. 1 for a graphical illustration of the system setup. The geometry is described in cylindrical coordinates $(r,\theta,z)$, where $r$ is the radial distance from the wedge apex, $\theta$ is the polar angle measured from the centerline between the walls, and $z$ is the axial direction along the wedge edge, which is invariant under translation. For $\alpha=\pi/4$, the wedge forms a right angle at the apex, whereas the planar wall limit is recovered for $\alpha=\pi/2$. No-slip boundary conditions, corresponding to zero fluid velocity, are imposed at $\theta=\pm\alpha$. The flow singularity, which may be either a point force or a point torque, is located at $(r,\theta,z)=(\rho,\beta,0)$.
At low Reynolds numbers, where the characteristic length and velocity scales of the flow are small, inertial effects become negligible compared with viscous forces. Under these conditions, the fluid motion is dominated by viscosity. Consequently, the governing equations reduce to the Stokes equations [72],
@@
The general solution of Eq. (1) can be expressed using the Papkovich–Neuber representation as [73, *Papkovich1932french, 75, 76]
@@
Note that $\Phi_{r}$ and $\Phi_{\theta}$ are not harmonic functions, a point that should be emphasized to avoid confusion in the subsequent developments.
In the present context, it is worth noting that in Stokes flow the solution is commonly represented in terms of three independent harmonic functions, known as Lamb’s general solution [78]. These are the harmonic velocity potential, the harmonic stream function, and the harmonic pressure field. Since the Papkovich–Neuber representation involves four harmonic functions, one of them is necessarily linearly dependent on the others. However, regardless of the specific choice made, the same physical solution is represented by this formulation.
@@
imposed at $\theta=\pm\alpha$. Accordingly, we require $\Pi=0$ at $\theta=\pm\alpha$, ensuring that $\partial\Pi/\partial r=\partial\Pi/\partial z=0$ at the boundaries.
@@
Our goal is to determine the complementary solution for a point force or point torque oriented along an arbitrary direction relative to the wedge.
@@ III.1 General overview
The FKL transform is a two-step integral transform commonly employed to simplify the analysis of problems with wedge geometries. In this approach, the Fourier transform is first applied along the axial coordinate, converting the dependence on this variable into a spectral representation. Subsequently, the KL transform is applied to the radial coordinate of the function, effectively mapping the radial dependence onto a continuous spectral parameter. This combined application of the Fourier and KL transforms reduces the original partial differential equations governing the system to a set of ordinary differential equations in the polar angle, greatly facilitating both analytical and numerical treatments.
In this review article, the hyperbolic sine, cosine, and tangent functions are denoted by sh, ch, and th, respectively. Additionally, sch and csh represent the hyperbolic secant and cosecant hyperbolic functions, respectively, defined as $\operatorname{sch}\zeta=1/\operatorname{ch}\zeta$ and $\operatorname{csh}\zeta=1/\operatorname{sh}\zeta$.
@@
$\hat{f}(r,k)=\frac{1}{\pi^{2}}\int_{0}^{\infty}\widetilde{f}(\nu,k)\,K_{i\nu}(|k|r)\operatorname{sh}(\pi\nu)\,\nu\,\mathrm{d}\nu\,.$ (18)
@@
$f(r,z)=\frac{1}{\pi}\int_{-\infty}^{\infty}\hat{f}(r,k)\,e^{-ikz}\,\mathrm{d}k\,.$ (19)
@@
We note that, as $\zeta\to\infty$, the modified Bessel function of the second kind defined by Eq. (15) exhibits the following asymptotic behavior [17, 16]
@@ III.2.1 Functional relationships
Here, we summarize key properties of the FKL transform that are relevant for solving Stokes flow problems involving a point force or point torque singularity. These include the FKL transform of derivatives with respect to $r$ and $z$, operations such as division by $r$ and multiplication by $z/r$, as well as the FKL transform of the Laplace equation. Proofs of these properties are provided in the Appendix of Ref. 56, so they are not repeated here. The results are presented in Tab. 1 to facilitate easy reference for the reader.
It should be noted that when the transformed function $f$ is an even function of $z$, as is the case here, its FKL transform is real. Consequently, for transforms that preserve $z$-parity, the FKL transform remains real, and the two terms defining the transform of $\partial f/\partial r$ and $f/r$ are complex conjugates, allowing the transform to be obtained by taking twice the real part of either term. In contrast, for transforms that reverse $z$-parity, the transform of an even function is purely imaginary. In this case, the two terms of the transforms of $\partial f/\partial z$ and $(z/r)f$ are negative complex conjugates, with identical imaginary parts and real parts equal in magnitude but opposite in sign. We chose to present the results as shown in Tab. 1, as this allows for a more streamlined derivation when applying the properties of the FKL transform.
@@ III.2.2 FKL transform of $1/s$
In the context of three-dimensional Stokes flow induced by a singularity, the FKL transform of $1/s$ is a fundamental quantity that helps construct the free-space contribution to the solution. For a point force, the harmonic functions are expressed directly in terms of $1/s$, whereas for a point torque they are given in terms of partial derivatives of $1/s$.
@@
As outlined above, the determination of the FKL transform proceeds by first evaluating the Fourier transform, followed by the KL transform. We seek the Fourier transform of $1/s$ as defined by Eq. (13). Taking into account the parity with respect to $z$, we have
@@
Note that for general complex numbers $a$ and $b$, the integral given in Eq. (27) is defined when $|\operatorname{Re}a|+|\operatorname{arg}b|<\pi$.
@@
This result constitutes a central element of the FKL method when applied to three-dimensional Stokes flow near wedges. Note that $a\in(0,\pi]\,.$
@@
Table 2: Nonzero coefficients defining $A_{j}$ and $A_{j}^{\dagger}$ for various orientations of point forces and torques; c.f. Eq. (32). Note that a transverse force and a parallel torque exhibit the same pattern of non-vanishing coefficients, as required.
@@ III.3.1 Representation of the solution in FKL space
The purpose of this general form is to significantly facilitate the inverse Fourier transform when obtaining the solution in real space. Table 2 lists the non-vanishing coefficients defining $A_{j}$ and $A_{j}^{\dagger}$ for an axial point force $F_{\parallel}$, a transverse point force $F_{\perp}$, an axial point torque $T_{\parallel}$, and a transverse point torque $T_{\perp}$. Note that a transverse force and a parallel torque share the same set of non-vanishing coefficients, as expected.
For the transverse force, we define $\delta$ as the angle between the force direction and the radial axis, i.e., $\delta=\tan^{-1}(F_{r}/F_{\theta})$, where $F_{r}$ and $F_{\theta}$ denote the radial and azimuthal components of the transverse force, respectively. Accordingly, $\delta=0$ corresponding to a purely radial force and $\delta=\pi/2$ to a purely azimuthal force. The same definition applies to the angle for a transverse torque normal to the wedge edge. Thanks to the linearity of the governing equations, any general motion can be described by a linear superposition of the solutions for axial and transverse singularities.

### 09 Apr 2026 v2

@@ I.1 The origin of the Kontorovich–Lebedev Transform
The KL integral transform is a relatively less known yet powerful integral transform, particularly well suited for solving boundary-value problems in wedge-shaped or conical geometries. The method was first developed in the late 1930s by the Russian mathematicians Mikhail Iosifovich Kontorovich (1906–1987) and Nikolai Nikolaevich Lebedev (1911–1994), who were at that time based in Leningrad, now Saint Petersburg, Russia. Using this approach, boundary-value problems in diffraction theory [7, 8] and electrodynamics [9] were solved by Kontorovich and Lebedev. The foundational mathematical theory of the integral transform was subsequently developed by Lebedev [10, *lebedev1946french, 12, 13]. A concise textbook presenting detailed solutions to applied mathematics problems—including applications of KL transforms—was coauthored by Lebedev together with Skal’skaia and Uflyand, first published in Russian and later translated and edited by Silverman [14, Ch. 6, p. 194].
In contrast to more familiar transforms, the KL transform is classified as an index integral transform [15, 16, 6]. These form a class of integral operators in which the transformed function depends on an index parameter, typically linked to special functions [17]. Over recent years, numerous significant contributions to the foundations of the KL transform have been made by Yakubovich and his collaborators [18, 19, 20, 21, 22, 23]. They advanced the theory of the KL transform by developing related integral and convolution operators and establishing connections with other integral transforms. Together, these developments established the KL transform as a versatile and rigorous tool for tackling complex boundary-value problems in mathematical physics. For a compact overview of this outstanding integral transform, including tables of transforms for common special functions, see the classic textbook by Erdélyi et al. [1, Ch. XII, p. 173].
@@ I.3 Applications in linear elasticity
Unlike its extensive use in electromagnetic and acoustic diffraction, the KL transform has seen comparatively limited application in linear elasticity [48]. The integral transform has been used to derive analytical solutions of the Navier–Cauchy equations for wedge-like geometries, providing a powerful tool for addressing stress and displacement fields in solids with corners. Early applications by Kassir and collaborators [49, 50] focused on solving crack problems using Papkovich–Neuber representations, yielding the stress distribution within the solid. To understand crack propagation and stress concentration in layered wedge-shaped structures, a periodic crack problem in an elastic layered wedge was analyzed by Pozharskii et al. [51]. In a related study, strip and elliptical cuts located in the middle of an elastic wedge were investigated by Aleksandrov and Pozharskii [52].
@@ I.4 Applications in low-Reynolds-number flows
The study of low-Reynolds-number flows near three-dimensional corners has provided fundamental insights into viscous fluid behavior in corners. Stokes flows in wedge geometries were first studied by Hasimoto and collaborators, who analyzed a sequence of wedge geometries, starting with a right-angled wedge [59], followed by obtuse wedges [60, 61, 62, 63], and semi-infinite planes [64]. The case of salient wedges was later investigated by Kim [65]. In these problems, the solution is obtained using a Fourier transform along the axial direction and a KL transform along the radial direction. These studies enabled the calculation of hydrodynamic mobility functions of colloidal particles in the far-field limit. Remarkably, it has been shown that when the wedge opening angle falls below a threshold value, an infinite sequence of eddy rings forms around the direction of the point force [66]. The Stokes flow in a semi-infinite wedge was later investigated by Shankar using the method of vector eigenfunctions [67].
@@
Although the KL transform was historically developed in wave physics, the present review concentrates on its role in viscous flows near wedges and corners. We revisit key techniques introduced in earlier works and summarize the solution methodology as adapted to incompressible Stokes flow. We focus on the steady Stokes equations, allowing the analysis to concentrate on spatial variations of the flow within wedge geometries. Since our approach combines Fourier transforms along the axial direction with the KL transform, we refer to this as the Fourier–Kontorovich–Lebedev (FKL) transform, representing a synthesis of both techniques.
Our discussion is restricted to systems in which the fluid motion is driven by either a point force or a point torque, and we focus on the physically relevant case of no-slip boundary conditions imposed on the wedge surfaces. Other types of boundary conditions, such as the free-slip condition, can also be analyzed within the same methodological framework presented in this article [57, 58]. The purpose of this review is to progressively develop the theoretical approach underlying the FKL transform, clarify its practical implementation for wedge geometries, and consolidate existing results to provide a coherent foundation for future advances in corner and wedge hydrodynamics. This review serves as an introductory read for those wishing to explore the modeling of flow problems near wedges and corners, progressively presenting the underlying mathematical framework and methods.
The remainder of the article is organized as follows. In Sec. II, we present the equations governing fluid motion and outline the framework used to obtain solutions based on the Papkovich–Neuber representation. Section III introduces the FKL transform, first by reviewing the relevant properties, and then by representing the solution form in FKL space. In Sec. IV, we summarize previously known solutions for a point force and point torque acting in a wedge-shaped geometry, oriented either parallel or perpendicular to the edge of the wedge. Finally, Sec. V provides a discussion and concluding remarks.
@@ II Governing equations
The system under investigation is three dimensional and consists of a viscous, incompressible fluid confined within a wedge-shaped domain, bounded by two walls located at $\theta=\pm\alpha$. The upper and lower walls of the wedge are characterized by the normal unit vectors $\left(\sin\alpha,-\cos\alpha,0\right)^{\top}$ and $\left(\sin\alpha,\cos\alpha,0\right)^{\top}$, respectively, both pointing toward the fluid domain. The intersection of the two walls forms a corner with a wedge semi-opening angle $\alpha\in(0,\pi/2]$; see Fig. 1 for a graphical illustration of the system setup. The geometry is described in cylindrical coordinates $(r,\theta,z)$, where $r$ is the radial distance from the wedge apex, $\theta$ is the polar angle measured from the centerline between the walls, and $z$ is the axial direction along the wedge edge, which is invariant under translation. For $\alpha=\pi/4$, the wedge forms a right angle at the apex, whereas the planar wall limit is recovered for $\alpha=\pi/2$. No-slip boundary conditions, corresponding to zero fluid velocity, are imposed at $\theta=\pm\alpha$.
We refer to a singularity as an idealized, localized force or torque acting at a single point in the fluid, represented mathematically by a Dirac delta function. A point force singularity, also known as a Stokeslet, corresponds to a concentrated force applied at a point, generating a flow field that decays away from that location but is formally infinite at the point itself. Similarly, a point torque singularity, also known as a rotlet, represents a localized torque acting at a point, inducing a rotational flow field that is likewise singular at the point of application. Such singularities constitute fundamental solutions of the Stokes equations and serve as building blocks for more complex flow configurations. In the present work, we assume that the flow singularity, whether a point force or a point torque, is located in cylindrical coordinates at $(r,\theta,z)=(\rho,\beta,0)$.
At low Reynolds numbers, where the characteristic length and velocity scales of the flow are small, inertial effects become negligible compared with viscous forces. Under these conditions, the fluid motion is governed by the Stokes equations [73],
@@
The general solution of Eq. (1) can be expressed using the Papkovich–Neuber representation as [74, 75, 76]
@@
The functions $\Phi_{r}$ and $\Phi_{\theta}$ are not harmonic, a point that should be emphasized to avoid confusion in the subsequent developments.
In Stokes flow, the solution is commonly represented in terms of three independent harmonic functions, known as Lamb’s general solution [78]. These are the harmonic velocity potential, the harmonic stream function, and the harmonic pressure field. Since the Papkovich–Neuber representation involves four harmonic functions, one of them is necessarily linearly dependent on the others. However, regardless of the specific choice made, the same physical solution is represented by this formulation.
@@
imposed at $\theta=\pm\alpha$. Accordingly, we require $\Pi=0$ at $\theta=\pm\alpha$, ensuring that $\partial\Pi/\partial r=\partial\Pi/\partial z=0$ at the boundaries. This is merely a choice and is not unique, as discussed above, since there is flexibility in selecting the fourth dependent harmonic function. Because this choice greatly simplifies the equations for the radial and axial components, setting $\Pi=0$ at the boundary is the standard approach when studying wedge problems using the Papkovich–Neuber representation.
@@ III.1 General overview
The FKL transform is a two-step integral transform used to simplify the analysis of problems involving wedge geometries. In this approach, the Fourier transform is first applied along the axial coordinate, converting the dependence on this variable into a spectral representation. Subsequently, the KL transform is applied to the radial coordinate of the function, effectively mapping the radial dependence onto a continuous spectral parameter. This combined application of the Fourier and KL transforms reduces the original partial differential equations governing the system to a set of ordinary differential equations in the polar angle, greatly facilitating both analytical and numerical treatments.
In this review article, the hyperbolic sine, cosine, tangent, and cotangent functions are denoted by sh, ch, th, and cth respectively. Additionally, sch and csh represent the hyperbolic secant and cosecant hyperbolic functions, respectively, defined as $\operatorname{sch}\zeta=1/\operatorname{ch}\zeta$ and $\operatorname{csh}\zeta=1/\operatorname{sh}\zeta$.
@@
$\hat{f}(r,k)=\frac{2}{\pi^{2}}\int_{0}^{\infty}\widetilde{f}(\nu,k)\,K_{i\nu}(|k|r)\operatorname{sh}(\pi\nu)\,\nu\,\mathrm{d}\nu\,.$ (18)
@@
$f(r,z)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\hat{f}(r,k)\,e^{-ikz}\,\mathrm{d}k\,.$ (19)
@@
As $\zeta\to\infty$, the modified Bessel function of the second kind defined by Eq. (15) exhibits the following asymptotic behavior [17, 16]
@@ III.2.1 Functional relationships
Here, we summarize key properties of the FKL transform that are relevant for solving Stokes flow problems involving a point force or point torque singularity. These include the FKL transform of derivatives with respect to $r$ and $z$, operations such as division by $r$ and multiplication by $z/r$, as well as the FKL transform of the Laplace equation. Proofs of these properties are provided in the Appendix of Ref. 57, so they are not repeated here. The results are presented in Tab. 1 to facilitate easy reference for the reader.
When the transformed function $f$ is an even function of $z$, as is the case here, its FKL transform is real. Consequently, for transforms that preserve $z$-parity, the FKL transform remains real, and the two terms defining the transform of $\partial f/\partial r$ and $f/r$ are complex conjugates, allowing the transform to be obtained by taking twice the real part of either term. In contrast, for transforms that reverse $z$-parity, the transform of an even function is purely imaginary. In this case, the two terms of the transforms of $\partial f/\partial z$ and $(z/r)f$ are negative complex conjugates, with identical imaginary parts and real parts equal in magnitude but opposite in sign. We chose to present the results as shown in Tab. 1, as this allows for a more streamlined derivation when applying the properties of the FKL transform.
@@ III.2.2 FKL transform of $1/s$
In the context of three-dimensional Stokes flow induced by a force or torque singularity, the FKL transform of $1/s$ is a fundamental quantity that helps construct the free-space contribution to the solution. For a point force, the harmonic functions are expressed directly in terms of $1/s$, whereas for a point torque they are given in terms of partial derivatives of $1/s$.
@@
As outlined above, the determination of the FKL transform proceeds by first evaluating the Fourier transform, followed by the KL transform. We determine the Fourier transform of $1/s$ as defined by Eq. (13). Taking into account the parity with respect to $z$, we have
@@
For general complex numbers $a$ and $b$, the integral given in Eq. (27) is defined when $|\operatorname{Re}a|+|\operatorname{arg}b|<\pi$.
@@
This result constitutes a central element of the FKL method when applied to three-dimensional Stokes flow near wedges. Here, $a\in(0,\pi]\,.$
@@
Table 2: Nonzero coefficients defining $A_{j}$ and $A_{j}^{\dagger}$ for various orientations of point forces and torques; c.f. Eq. (32). A transverse force and a parallel torque exhibit the same pattern of non-vanishing coefficients, as required.
@@ III.3.1 Representation of the solution in FKL space
The purpose of this general form is to significantly facilitate the inverse Fourier transform when obtaining the solution in real space. Table 2 lists the non-vanishing coefficients defining $A_{j}$ and $A_{j}^{\dagger}$ for an axial point force $F_{\parallel}$, a transverse point force $F_{\perp}$, an axial point torque $T_{\parallel}$, and a transverse point torque $T_{\perp}$. A transverse force and a parallel torque share the same set of non-vanishing coefficients, as expected.
For the transverse force, we define $\sigma$ as the angle between the force direction and the radial axis, i.e., $\sigma=\tan^{-1}(F_{r}/F_{\theta})$, where $F_{r}$ and $F_{\theta}$ denote the radial and azimuthal components of the transverse force, respectively. Accordingly, $\sigma=0$ corresponding to a purely radial force and $\sigma=\pi/2$ to a purely azimuthal force. The same definition applies to the angle for a transverse torque normal to the wedge edge. Thanks to the linearity of the governing equations, any general motion can be described by a linear superposition of the solutions for axial and transverse singularities.